An ore contains Fe3O4 and no other iron. The iron in a 56.7-gram sample of the ore is all converted by a series of chemical reactions to Fe2O3. The mass of Fe2O3 is measured to be 26.7 g. What was the mass of Fe3O4 in the sample of ore?
1
Expert's answer
2019-08-27T04:44:59-0400

"Fe_3O_4" is converted to "Fe_2O_3" by the following reaction:

"4Fe_3O_4+O_2 \to 6 Fe_2O_3"

As per the stoichiometry of the reaction,

6 moles of "Fe_2O_3" is equivalent to 4 moles of "Fe_3O_4"

or, "6 \times (56 \times 2 +16 \times 3) \space gm \space Fe_2O_3 \equiv"

"4 \times(56 \times 3 +16 \times 4)\space gm \space Fe_3O_4"

or,"6 \times 160 \space gm Fe_2O_3 \equiv 4 \times 232 \space gm \space Fe_3O_4"

or,"26.7 \space gm \space Fe_2O_3\equiv\frac{4 \times 232}{6 \times 160}\times26.7 \space gm \space Fe_3O_4"

or, 26.7 grams of "Fe_2O_3" will have 25.81 grams of "Fe_3O_4" in the ore sample.